# A Universal Knowledge Embedded Contrastive Learning Framework for Hyperspectral Image Classification

### v1

@@ Abstract
Hyperspectral image (HSI) classification techniques have been intensively studied and a variety of models have been developed. However, these HSI classification models are confined to pocket models and unrealistic ways of datasets partitioning. The former limits the generalization performance of the model and the latter is partitioned leads to inflated model evaluation metrics, which results in plummeting model performance in the real world. Therefore, we propose a universal knowledge embedded contrastive learning framework (KnowCL) for supervised, unsupervised, and semisupervised HSI classification, which largely closes the gap of HSI classification models between pocket models and standard vision backbones. We present a new HSI processing pipeline in conjunction with a range of data transformation and augmentation techniques that provide diverse data representations and realistic data partitioning. The proposed framework based on this pipeline is compatible with all kinds of backbones and can fully exploit labeled and unlabeled samples with expected training time. Furthermore, we design a new loss function, which can adaptively fuse the supervised loss and unsupervised loss, enhancing the learning performance. This proposed new classification paradigm shows great potentials in exploring for HSI classification technology. The code can be accessed at https://github.com/quanweiliu/KnowCL.
@@ I Introduction
To fully exploit the information contained in HSIs, previous work focus on machine learning models that first extract a series of features of HSIs, such as spectral features, texture features, morphological semantic features, etc. Then the extracted feaures are fed into classifiers such as $k$-nearest neighbor classifier (KNN) to obtain classification results[6, 7, 8]. However, early machine learning methods relied on handcrafted shallow features and empirical assumptions that seriously hinder the model building efficiency and generalization performance.
@@
In addition to the learning strategies, another critical step is the evaluation of HSI classification methods, such as benchmark datasets, sampling strategies and metrics, where sampling strategies have become one of the most urgent methods to be revised [21]. A common sampling setting in the HSI classification is to randomly select a fixed number or a fixed proportion of labeled samples from an HSI as a training set and the rest of the samples in this image as a test set [22]. These selected samples are randomly distributed throughout the image. The spatial autocorrelation of the land cover makes similar features share the same label with a high probability [23]. Consequently, the random sampling based on pixel-by-pixel is very inefficient and impractical. In a real environment, most regions rather than pixels are usually identified as ground truth by visual interpretation or field survey. On the other hand, with the development of HSI classification based on CNNs, spectral-spatial classification becomes a concise and efficient method. However, the classification accuracy is appreciably high due to the spatial autocorrelation in the test data set. A considerable number of work has achieved nearly 100% correctness in conditions where only a small number of training samples are used [24, 25, 26, 27]. Cao et. al. [28] conducted a preliminary performance comparison of HSI classification based on the overlapped and nonoverlapped sampling strategy, where the experimental results also showed that the classification accuracy is significantly reduction using the nonoverlapped sampling strategy.
@@
Given the problems with modeling and evaluation, in this article, we propose a knowledge-embedded contrastive learning framework (KnowCL) for supervised, unsupervised, and semi-supervised HSI classification. The framework diagram is shown in Fig. 2. First, we adopt a disjoint approach to divide the HSIs into training and test sets, and this sampling strategy more realistically reproduces the labeling of HSIs. Then, we obtain a diversity of labeled and unlabeled samples through a novel HSI processing pipeline. The supervised classification model is constructed with an iteration optimization of a backbone and a supervised head. The unsupervised model is optimized to maximize the similarity of the two feature extraction branches. We introduced task-specific knowledge into the unsupervised model to construct a novel semi-supervised learning model, which will greatly speed up the network fitting and enhance the model performance. Moreover, the auto-weighted loss that can adaptively fuse the the supervised learning and contrastive learning is designed for the framework. In addition, we present a light vision transformer (ViT) and test ResNet as backbones to validate the effectiveness of our framework. The main contributions of this research are listed as follows.
@@ II-B Unsupervised Learning
Contrastive learning centers on making positive sample pairs closer to each other and negative sample pairs farther apart [43, 44]. For the image, sample pairs can be constructed using image augmentation methods (e.g., rotation, flip, blur, etc.). A sample undergoing a series of data augmentations produces two samples as a pair of positive samples, and other samples from the same batch are viewed as negative sample pairs. MoCo [39] builds a sequence of stored negative samples to expand the number of negative samples. SimCLR [42] uses 128 TPUs directly to train the model to store a large batch size samples. DMVL[45] uses the infoNCE loss function similar to the SimCLR framework to construct the HSI classification model. XDCL[27] learns domain invariance features in the spectral and spatial domains of HSIs.
@@ III-A Data Processing
CL requires learning invariant feature representations of objects from different views of the same object. Data augmentation is widely used for CL as a simple way to generate multi-view representations. HSIs have multiple spectral bands. Beyond conventional data augmentation, these bands can all be viewed as different observations of the same target. We halve the HSI into two groups for extracting invariant feature representation, as shown in Fig. 3. Considering the impact of different sensors and bands on the classification performance, we do not directly reduce each HSI to three dimensions to apply natural image augmentation methods. Instead, we use principal component analysis (PCA) [20] to extract the most discriminative features in each HSI group. With each grouped spectral pixel as the center, image blocks $X\in\mathbb{R}^{H\times{W}\times C}$ are constructed. The common data augmentation includes random crop and resize, random vertical and horizontal flips, and Gaussian blur. It is worth noting that HSIs do not have the concept of color, so data augmentation, such as color jitter, is not adaptive. The input HSI blocks $X$ are divided into two views after data processing: a labeled sample view $x$ and an unlabeled sample view $\hat{x}$.
@@ III-C Loss Functions and Fusion Strategy
${\cal{L}_{\cal T}}=\sum\limits_{\cal{t}\in T}{\frac{1}{{2\cdot w_{t}^{2}}}\cdot{L_{t}}}+\ln(1+w_{t}^{2}),$ (6)
@@ IV-A Reference Datasets
Four HSI datasets are mainly used in this study which are University of Pavia dataset (UP), Salinas dataset, Dioni dataset and 2018 IEEE GRSS data fusion contest dataset (DFC2018). These four datasets used four sensors and captured four regions at different times, as shown in Fig. 5. The DFC2018 has a minimum of 48 bands, while the Salinas has up to 204 bands. The DFC2018 and UP only scan the visible near-infrared spectra. The Salinas and Dioni acquire both visible near-infrared and shortwave infrared spectra. Dioni is located in the middle of the wilderness and has a low spatial resolution, while the remaining three datasets focus on urban and agricultural scenes and have a high spatial resolution. Of these four datasets, the Salinas was collected the earliest, which is relatively small in size. DFC2018 is approximately 12 times larger than the Salinas data, and also contains the largest number of labeled samples. The detailed parameters associated with these data are given in Table I.
As stated in the section III-A, we adopted a well disjoint data sampling strategy for UP, Salinas, Dioni datasets. For the DFC2018, the training and test set had been defined by the GRSS. The detailed division is shown in Table II and Fig. 5 visualizes the distribution of labels in the training set and test set.
@@ IV-B2 Technical details
This framework unifies supervised learning, unsupervised learning and semi-supervised learning, and we name them KnowCL-SU, KnowCL-US and knowCL, respectively. KnowCL(under kNN classifier) has the best performance and is mainly used to analyse and discuss the superiority of the proposed algorithm. We adopt a modified ViT as the backbone whose has four attention blocks. We first intercept HSI blocks of size $C\times 25\times 25$ based on each center pixel, where $C$ depends on the input HSI. Following the standard ViT, the image blocks are further flattened with patches size of 4. The input HSI blocks will be encoded into 126-D latent features. This features will be further mapped to 256-D latent features by a 1-layer MLP contrastive head and a 2-layer MLP supervisory head to compute the infoNCE loss and cross-entropy loss, respectively. The temperature $\tau$ is set as 0.5. The data augmentation is available in PyTorch’s torchvision package. The random crop size is set to 23×23-pixel, considering the experimental results. AdamW is used as the optimizer with a cosine annealing learning rate schedule. We train for 30 epochs with batch size 256, weight decay 0.000001, and learning rate 0.001. We follow [51, 52] to evaluate the performance of the model with a weighted kNN classifier, and the $k$ value is set to 5 according to the performance of the experiment. We also further tested our model on the supervised head and linear classification protocol to explore the accuracy limits of the model. All models are implemented based on Pytorch 1.9 and performed with an NVIDIA GeForce RTX 3090 GPU.
@@ IV-C2 Batch size
Table IV shows the impact of batch size when models are trained on the DFC2018. Unlike simCLR [42], model performance does not improve with increasing batch size. On the one hand, it may be due to the fact that the training data involved in the training of the model are fewer, and they are taken from several regions with more similar features so global features can be learned with a smaller batch size. On the other hand, it suggests that the insertion of labeling knowledge allows the model to learn task-specific knowledge, giving the model enough power to find the optimal feature representation under a small batch size.
@@ IV-C3 Bands
HSIs contain bands far beyond those of natural images. The previous work either directly input all the bands or took only 3 bands without considering the effect of different sensors and feature types on the classification performance. However, the spectral correlation affects its classification performance. PCA, a simple dimensionality reduction method is applied to remove the noise interference and extract the most discriminative feature. Table V shows the impact of the number of principal components when the model are trained on the four datasets. We find that when the number of principal components increases, the performance of the model will have a great difference. The overall presentation is an inverse U-shape. The results manifest that different hyperspectral sensors and scenes extract different principal components to achieve the best classification performance.
@@ IV-D Quantitative Comparison
For supervised and unsupervised learning, KnowCL-SU and KnowCL-US achieve performance levels that are on par with or surpassing comparison methods, while semi-supervised learning KnowCL combines transformer and CL to achieve the best accuracy with balancing time and computation. Moreover, KnowCL is able to extract more fine-grained feature representations than hyperpixel-based graph learning models. Compared to the unsupervised learning model, KnowCL is able to learn task-specific feature representations using label knowledge. In addition, the performance of the linear evaluation is slightly lower than KNN classifier, indicating that our model has fully utilized the label information in the training phase. As shown in Table VII, KnowCL also outperforms the other methods in all quality indices in Salinas. Salinas contains several crops that are uniformly distributed and easily distinguishable, and the other compared algorithms also achieve better classification results.
@@
Table VIII and IX show the experimental results on the Dioni and DFC2018. It is worth noting that the memory requirement of traditional semi-supervised learning methods increases exponentially with the amount of data. Conventional GPU are no longer sufficient to support the operation of Dioni and DFC2018. Therefore, only supervised and unsupervised learning methods are used as comparison algorithms. Larger HSIs contain more diverse feature classes and produce more complex spectral combinations, leading to the accuracy of these two datasets having a large gap compared to UP and Salinas. It is in line with the performance of these algorithms when applied to the real world. We can also observe that the linear probe results are higher than the kNN classification results on the DFC2018. This can also be attributed to the complexity of the feature components of DFC2018. The large number of mixed pixels [54] leads to poor classification performance. The best results obtained with KnowCL can substantially outperform comparative models.
@@ IV-F Ablation Study
Fig. 12 visualizes the feature maps using the supervised learning, unsupervised learning, additive fusion learning and adaptive fusion learning strategies, respectively. We selectively pick out some representative feature maps for visual comparison, where the visualization results of using the ${\cal L}_{cl}+{\cal L}_{ce}$ and ${\cal L}_{\cal T}$ learning strategies have finer appearance than those using only one loss function. For example, the activation maps for the region in the white box show that the non-residential buildings area based on the fusion loss has sharper edges, object contours, and texture structures. This also visually confirms the superiority of the designed framework.
@@ IV-G Backbones and classifier
There are numerous models for HSI classification at present. We try to propose a general framework which is compatible with the mainstream deep learning backbone. Table X shows the performance of three ViT-based networks and three ResNet-based networks as well as their number of parameters under kNN, supervised head and linear classification protocols. Fig. 1 provides a more visual comparison of them with the other models in DFC2018. We find that the ViT-based networks have similar classification performance under the three classification strategies, while the accuracy of the ResNet-based networks becomes higher at linear evaluation as the model parameters increase. We surmise that this is because ViT is able to extract global features at each layer, so the model performance converges faster. ResNet has the local receptive fields and requires deeper structure to adequately extract global features. Moreover, our model achieves a large advantage in all parametric quantities.
@@ V Conclusion
In this paper, we presents a universal HSI classification framework designed for addressing real-world HSI classification problems, which is compatible with supervised learning, unsupervised learning and semi-supervised learning. Among them, semi-supervised learning classification KnowCL is newly defined to fit the requirements of real applications. An efficient contrastive learning strategy is constructed by means of labeled knowledge embedding, where the parameters, FLOPs and execution time is comparable to that of supervised learning but can fully extract features from unlabeled samples. We conduct four disjoint datasets, and a large number of experiments on them quantitatively and intuitively validate the superiority of the proposed model. Future, it is possible to adopt pre-trained visual foundation model for HSI classification, which will greatly advance the application of HSIs in the real world.

### v3

@@ Abstract
Hyperspectral image (HSI) classification techniques have been intensively studied and a variety of models have been developed. However, these HSI classification models are confined to pocket models and unrealistic ways of dataset partitioning. The former limits the generalization performance of the model and the latter is partitioned leading to inflated model evaluation metrics, which results in plummeting model performance in the real world. Therefore, we propose a universal knowledge embedded contrastive learning framework (KnowCL) for supervised, unsupervised, and semisupervised HSI classification, which largely closes the gap between HSI classification models between pocket models and standard vision backbones. We present a new HSI processing pipeline in conjunction with a range of data transformation and augmentation techniques that provide diverse data representations and realistic data partitioning. The proposed framework based on this pipeline is compatible with all kinds of backbones and can fully exploit labeled and unlabeled samples with the expected training time. Furthermore, we design a new loss function, which can adaptively fuse the supervised loss and unsupervised loss, enhancing the learning performance. This proposed new classification paradigm shows great potential in exploring for HSI classification technology. The code can be accessed at https://github.com/quanweiliu/KnowCL.
@@ I Introduction
To fully exploit the information contained in HSIs, previous work focuses on machine learning models that first extract a series of features of HSIs, such as spectral features, texture features, morphological semantic features, etc. Then the extracted features are fed into classifiers such as $k$-nearest neighbor classifier (KNN) to obtain classification results[6, 7, 8]. However, early machine learning methods relied on handcrafted shallow features and empirical assumptions that seriously hindered the model-building efficiency and generalization performance.
@@
In addition to the learning strategies, another critical step is the evaluation of HSI classification methods, such as benchmark datasets, sampling strategies and metrics, where sampling strategies have become one of the most urgent methods to be revised [21]. A common sampling setting in the HSI classification is to randomly select a fixed number or a fixed proportion of labeled samples from an HSI as a training set and the rest of the samples in this image as a test set [22]. These selected samples are randomly distributed throughout the image. The spatial autocorrelation of the land cover makes similar features share the same label with a high probability [23]. Consequently, the random sampling based on pixel-by-pixel is very inefficient and impractical. In a real environment, most regions rather than pixels are usually identified as ground truth by visual interpretation or field survey. On the other hand, with the development of HSI classification based on CNNs, spectral-spatial classification becomes a concise and efficient method. However, the classification accuracy is appreciably high due to the spatial autocorrelation in the test data set. A considerable number of studies have achieved nearly 100% correctness in conditions where only a small number of training samples are used [24, 25, 26, 27]. Cao et. al. [28] conducted a preliminary performance comparison of HSI classification based on the overlapped and nonoverlapped sampling strategy, where the experimental results also showed that the classification accuracy is significantly reduction using the nonoverlapped sampling strategy.
@@
Given the problems with modeling and evaluation, in this article, we propose a knowledge-embedded contrastive learning framework (KnowCL) for supervised, unsupervised, and semi-supervised HSI classification. The framework diagram is shown in Fig. 2. First, we adopt a disjoint approach to divide the HSIs into training and test sets, and this sampling strategy more realistically reproduces the labeling of HSIs. Then, we obtain a diversity of labeled and unlabeled samples through a novel HSI processing pipeline. The supervised classification model is constructed with an iteration optimization of a backbone and a supervised head. The unsupervised model is optimized to maximize the similarity of the two feature extraction branches. We introduced task-specific knowledge into the unsupervised model to construct a novel semi-supervised learning model, which will greatly speed up the network fitting and enhance the model performance. Moreover, the auto-weighted loss that can adaptively fuse supervised learning and contrastive learning is designed for the framework. In addition, we present a light vision transformer (ViT) and test ResNet as backbones to validate the effectiveness of our framework. The main contributions of this research are listed as follows.
@@ II-B Unsupervised Learning
Contrastive learning centers on making positive sample pairs closer to each other and negative sample pairs farther apart [43, 44]. For the image, sample pairs can be constructed using image augmentation methods (e.g., rotation, flip, blur, etc.). A sample undergoing a series of data augmentations produces two samples as a pair of positive samples, and other samples from the same batch are viewed as negative sample pairs. MoCo [39] builds a sequence of stored negative samples to expand the number of negative samples. SimCLR [42] uses 128 TPUs directly to train the model to store large batch-size samples. DMVL[45] uses the infoNCE loss function similar to the SimCLR framework to construct the HSI classification model. XDCL[27] learns domain invariance features in the spectral and spatial domains of HSIs.
@@ III-A Data Processing
CL requires learning invariant feature representations of objects from different views of the same object. Data augmentation is widely used for CL as a simple way to generate multi-view representations. HSIs have multiple spectral bands. Beyond conventional data augmentation, these bands can all be viewed as different observations of the same target. We halve the HSI into two groups for extracting invariant feature representation, as shown in Fig. 3. Considering the impact of different sensors and bands on the classification performance, we do not directly reduce each HSI to three dimensions to apply natural image augmentation methods. Instead, we use principal component analysis (PCA) [20] to extract the most discriminative features in each HSI group. With each grouped spectral pixel as the center, image blocks $X\in\mathbb{R}^{H\times{W}\times C}$ are constructed. The common data augmentation includes random crop and resize, random vertical and horizontal flips, as well as Gaussian blur. It is worth noting that HSIs do not have the concept of color, so data augmentation, such as color jitter, is not adaptive. The input of HSI blocks $X$ is divided into two views after data processing: a labeled sample view $x$ and an unlabeled sample view $\hat{x}$.
@@ III-C Loss Functions and Fusion Strategy
${\cal{L}_{\cal T}}=\sum\limits_{t\in T}{\frac{1}{{2\cdot w_{t}^{2}}}\cdot{L_{t}}}+\ln(1+w_{t}^{2}),$ (6)
@@ IV-A Reference Datasets
Four HSI datasets are mainly used in this study which are the University of Pavia dataset (UP), the Salinas dataset, the Dioni dataset and the 2018 IEEE GRSS data fusion contest dataset (DFC2018). These four datasets used four sensors and captured four regions at different times, as shown in Fig. 5. The DFC2018 has a minimum of 48 bands, while the Salinas has up to 204 bands. The DFC2018 and UP only scan the visible near-infrared spectra. The Salinas and Dioni acquire both visible near-infrared and shortwave infrared spectra. Dioni is located in the middle of the wilderness and has a low spatial resolution, while the remaining three datasets focus on urban and agricultural scenes and have a high spatial resolution. Of these four datasets, the Salinas was collected the earliest, which is relatively small in size. DFC2018 is approximately 12 times larger than the Salinas data and also contains the largest number of labeled samples. The detailed parameters associated with these data are given in Table I.
As stated in the section III-A, we adopted a well-disjoint data sampling strategy for UP, Salinas, and Dioni datasets. For the DFC2018, the training and test set had been defined by the GRSS. The detailed division is shown in Table II and Fig. 5 visualizes the distribution of labels in the training set and test set.
@@ IV-B2 Technical details
This framework unifies supervised learning, unsupervised learning and semi-supervised learning, and we name them KnowCL-SU, KnowCL-US and knowCL, respectively. KnowCL(under the kNN classifier) has the best performance and is mainly used to analyse and discuss the superiority of the proposed algorithm. We adopt a modified ViT as the backbone which has four attention blocks. We first intercept HSI blocks of size $C\times 25\times 25$ based on each center pixel, where $C$ depends on the input HSI. Following the standard ViT, the image blocks are further flattened with patches size of 4. The input HSI blocks will be encoded into 126-D latent features. These features will be further mapped to 256-D latent features by a 1-layer MLP contrastive head and a 2-layer MLP supervisory head to compute the infoNCE loss and cross-entropy loss, respectively. The temperature $\tau$ is set as 0.5. The data augmentation is available in PyTorch’s torchvision package. The random crop size is set to 23×23-pixel, considering the experimental results. AdamW is used as the optimizer with a cosine annealing learning rate schedule. We train for 30 epochs with batch size 256, weight decay 0.000001, and learning rate 0.001. We follow [51, 52] to evaluate the performance of the model with a weighted kNN classifier, and the $k$ value is set to 5 according to the performance of the experiment. We also further tested our model on the supervised head and linear classification protocol to explore the accuracy limits of the model. All models are implemented based on Pytorch 1.9 and performed with an NVIDIA GeForce RTX 3090 GPU.
@@ IV-C2 Batch size
Table IV shows the impact of batch size when models are trained on the DFC2018. Unlike the simCLR [42], model performance does not improve with increasing batch size. On the one hand, it may be due to the fact that the training data involved in the training of the model are fewer, and they are taken from several regions with more similar features so global features can be learned with a smaller batch size. On the other hand, it suggests that the insertion of labeling knowledge allows the model to learn task-specific knowledge, giving the model enough power to find the optimal feature representation under a small batch size.
@@ IV-C3 Bands
HSIs contain bands far beyond those of natural images. The previous work either directly input all the bands or took only 3 bands without considering the effect of different sensors and feature types on the classification performance. However, the spectral correlation affects its classification performance. PCA, a simple dimensionality reduction method is applied to remove the noise interference and extract the most discriminative feature. Table V shows the impact of the number of principal components when the model is trained on the four datasets. We find that when the number of principal components increases, the performance of the model will have a great difference. The overall presentation is an inverse U-shape. The results manifest that different hyperspectral sensors and scenes extract different principal components to achieve the best classification performance.
@@ IV-D Quantitative Comparison
For supervised and unsupervised learning, KnowCL-SU and KnowCL-US achieve performance levels that are on par with or surpassing comparison methods, while semi-supervised learning KnowCL combines transformer and CL to achieve the best accuracy with balancing time and computation. Moreover, KnowCL is able to extract more fine-grained feature representations than hyperpixel-based graph learning models. Compared to the unsupervised learning model, KnowCL is able to learn task-specific feature representations using label knowledge. In addition, the performance of the linear evaluation is slightly lower than the KNN classifier, indicating that our model has fully utilized the label information in the training phase. As shown in Table VII, KnowCL also outperforms the other methods in all quality indices in Salinas. Salinas contains several crops that are uniformly distributed and easily distinguishable, and the other compared algorithms also achieve better classification results.
@@
Table VIII and IX show the experimental results on the Dioni and DFC2018. It is worth noting that the memory requirement of traditional semi-supervised learning methods increases exponentially with the amount of data. Conventional GPUs are no longer sufficient to support the operation of Dioni and DFC2018. Therefore, only supervised and unsupervised learning methods are used as comparison algorithms. Larger HSIs contain more diverse feature classes and produce more complex spectral combinations, leading to the accuracy of these two datasets having a large gap compared to UP and Salinas. It is in line with the performance of these algorithms when applied to the real world. We can also observe that the linear probe results are higher than the kNN classification results on the DFC2018. This can also be attributed to the complexity of the feature components of DFC2018. The large number of mixed pixels [54] leads to poor classification performance. The best results obtained with KnowCL can substantially outperform comparative models.
@@ IV-F Ablation Study
Fig. 12 visualizes the feature maps using supervised learning, unsupervised learning, additive fusion learning and adaptive fusion learning strategies, respectively. We selectively pick out some representative feature maps for visual comparison, where the visualization results of using the ${\cal L}_{cl}+{\cal L}_{ce}$ and ${\cal L}_{\cal T}$ learning strategies have finer appearance than those using only one loss function. For example, the activation maps for the region in the white box show that the area of the non-residential building based on the fusion loss has sharper edges, object contours, and texture structures. This also visually confirms the superiority of the designed framework.
@@ IV-G Backbones and classifier
There are numerous models for HSI classification at present. We try to propose a general framework that is compatible with the mainstream deep learning backbone. Table X shows the performance of three ViT-based networks and three ResNet-based networks as well as their number of parameters under kNN, supervised head and linear classification protocols. Fig. 1 provides a more visual comparison of them with the other models in DFC2018. We find that the ViT-based networks have similar classification performance under the three classification strategies, while the accuracy of the ResNet-based networks becomes higher at linear evaluation as the model parameters increase. We surmise that this is because ViT is able to extract global features at each layer, so the model performance converges faster. ResNet has the local receptive fields and requires a deeper structure to adequately extract global features. Moreover, our model achieves a large advantage in all parametric quantities.
@@ V Conclusion
In this paper, we present a universal HSI classification framework designed for addressing real-world HSI classification problems, which is compatible with supervised learning, unsupervised learning and semi-supervised learning. Among them, the semi-supervised learning classification KnowCL is newly defined to fit the requirements of real applications. An efficient contrastive learning strategy is constructed by means of labeled knowledge embedding, where the parameters, FLOPs and execution time are comparable to that of supervised learning but can fully extract features from unlabeled samples. We conduct four disjoint datasets, and a large number of experiments on them quantitatively and intuitively validate the superiority of the proposed model. Future, it is possible to adopt a pre-trained visual foundation model for HSI classification, which will greatly advance the application of HSIs in the real world.